OpenStudy (anonymous):

Which of the following statements is true, given the following information? Density of A = 2.0 g/mL; Density of B = 3.5 g/mL; Density of C = 0.10 g/mL (All densities are at 200C)

OpenStudy (rushwr):

What are the statements

OpenStudy (rushwr):

????// @kheath39

OpenStudy (anonymous):

1 mL of A weighs more than 1 mL of B. 10 g of C has a smaller volume than 10 g of B. A, B and C all have densities greater than water at 20oC. One liter of B has a greater mass than 1 liter of C.

OpenStudy (rushwr):

What do u think ?

OpenStudy (rushwr):

1) wrong because 1ml of A weights 2g and 1mL of B weights 3.5g 2) D= m/v D= density V= volume m= mass So for C we know the density and the mass we can clearly find the volume. \[Volume =\frac{ mass }{ density } = \frac{ 10g }{ 0.10g} = 100 \] Volume of 10g of B = 10g/3.5 = 2.85 So that option is wrong

OpenStudy (rushwr):

3) is wrong because at 20^oC water has a density around 0.99gml^-1 C has a density lower than that. So wrong 4) Mass = density * volume Mass of B = 3.5gml^-1 * 1000ml = 3500g Mass of C = 0.1gml^-1 * 1000mL = 100g So mass of B > Mass of C so that option is correct.